This was from standardized test (ASSET):
Tractors A,B, C & D are equipped identically. Together A,B & C can plow 180 acres in 2 hrs. How many acres can all 4 (A, B, C & D) plow in 3 hrs?

Answer by ankor@dixie-net.com(22740)

Tractors A,B, C & D are equipped identically. Together A,B & C can plow 180 acres in 2 hrs. How many acres can all 4 (A, B, C & D) plow in 3 hrs?
:
Let T = acres plowed in 1 hr, by one tractor
3T = 180/2
6T = 180
T = 180/6
T = 30 acres per hr (1 tractor)
:
Using 4 tractors for 3 hrs
Acres = 4T * 3
Substitute 30 for T
Acres = 4(30) * 3
Acres = 120 * 3
Acres = 360 (by 4 tractors for 3 hrs.)
:
Did this make sense? HOpe so